OpenStudy (anonymous):

(I'LL BECOME A FAN AND GIVE YOU A MEDAL !! :)♥ I JUST REALLY NEED HELP PLEASE PLEASE) 1. What is the simplified form of the expression? 2x^2 + 4y+ 3x^2- 2y + 3y A. 5x^4 + 5y^2 B. 5x^2 + 5x^2y C. 5x^2 + 5y D. 10x^2y

OpenStudy (praxer):

just add the coefficients of x^2 and y respectively.. you will get it to be 5x^2+5y.. So, option C. is correct......... :)

OpenStudy (anonymous):

ohh okay. i got a few more questions :)

OpenStudy (anonymous):

c

OpenStudy (the_fizicx99):

Ask :>

OpenStudy (anonymous):

Simplify the following expression: 4(12x +16) A. 64x+48 B. 48x + 64 C. 12x + 64 D. 48x + 16

OpenStudy (anonymous):

b only multiply

OpenStudy (anonymous):

ohhh I see now. thankies ^_^ is it alright if I keep asking questions? heehee :c

OpenStudy (anonymous):

Which property is illustrated by the statement? 2(- 6/7) = -(6/7) 2 A. Associative Property of Addition B.Inverse Property of Multiplication C.Commutative Property of Multiplication D. Commutative Property of Addition

OpenStudy (praxer):

C.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

yay! thank you. I got a few more.. umm 10. What is the algebraic expression for the following word phrase: "the product of g and 4"? a. g-4 b. 4g c. g/4 d. g+4 YOU GUYS ARE LITERALLY SAVING ME RIGHT NOW♥♥

OpenStudy (praxer):

B. g*4= 4g

OpenStudy (the_fizicx99):

g * 4 = 4g

OpenStudy (anonymous):

why do they change my words? lmfao

OpenStudy (the_fizicx99):

You shouldn't curse, you can get suspended for that. I got suspended for it ;-;

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

I deleted my reply. am I safe?

OpenStudy (the_fizicx99):

Um moderators can see deleted replies. But if you weren't reported they won't bother. everything is monitored on here.

OpenStudy (anonymous):

okay. may I ask another question?

OpenStudy (the_fizicx99):

Make another post it starts to get laggy :/